# Quantifying Emergence in Large Language Models

### v1

@@ 3.2 How to Estimate Emergence in LLMs?
It is not feasible to directly compute the mutual information in Eq. 3 using Kullback-Leibler (KL) divergence, as the input lies in a high-dimensional continuous space. To address that, we resort to an approximation method using mean values proposed in [6]:
@@
where $\theta$ denotes the parameters for $f$, ||||| | denotes the concatenation operation and $B$ is the batch size. $x,y\in\mathbb{R}^{D}$ are two inputs for computing the mutual information $MI(x,y)$. $x^{b}||y^{b}$ corresponds to sampling from the joint distribution $P_{XY}$, while $x^{b}||y^{i\neq b}$ corresponds to sampling from the marginal distribution $P_{X}$ and $P_{Y}$ [^4] In our implementation, the batch size was increased to encompass the entirety of the sample set to ensure the rationality of $P_{XY}$, $P_{X}$ and $P_{Y}$.. When $\mathcal{L}$ converges to the minimum $\hat{\mathcal{L}}$, we can obtain the final estimated mutual information as $MI(x,y)$=$-\log_{2}^{e}*\hat{\mathcal{L}}$. (More details and proofs are shown in [6].)
@@
To get the emergence value $E(l)$ in Eq. 3, we compute $MI(h^{ma}_{l+1},h^{ma}_{l})$ by replacing $x^{b}$ and $y^{b}$ in Eq. 5 with $h^{ma}_{l+1,s}$ and $h^{ma}_{l,s}$ obtained by applying the LM to the same input sequence $s$, whereas replacing $y^{i\neq b}$ with $h^{ma}_{l,s^{\prime}}$ using a different sequence $s^{\prime}\neq s$. Similar operations apply when computing $MI(h^{mi\_t}_{l+1},h^{mi\_t}_{l})$. (Refer to Appendix A for a complete algorithm for estimating the emergence.)
3.3 Implementation
We examine the emergence strength of LLMs under two distinct settings: ICL with few-shot examples and natural sentences without demonstrations. For both settings, our algorithm requires that the number of samples is sufficiently large (over 300k) to provide a good estimate of the mutual information. Meanwhile, the length of each sequence within a dataset should be kept the same to facilitate position-wise observation and meaningful computation.
@@ 3.3 Implementation
Due to resource constraints, our comparative analysis is limited to GPT2-large (812M)  [^5] https://huggingface.co/openai-community/gpt2-large, GPT2-XL (1.61B), GEMMA (2.51B), and OpenLlama (3B)  [^6] https://huggingface.co/openlm-research/open_llama_3b models. Fortunately, this parameter range is sufficient to observe variations and regularities of the emergence phenomenon. For those LLMs with extremely large size or closed resource (e.g., GPT-4, Claude3, etc.), we design another strategy that enables their emergence evaluations as shown in Section 5.2. All computational experiments can be conducted on one NVIDIA GeForce RTX 3090 GPU [^7] For easy deployment, we split the representations of LLMs into several 768-dimension segments and the final representation is the mean of these segments.. The estimator $f$ in Eq. 3 is a model of a 10-layer neural network comprising linear layers and leaky ReLU activation functions, where each linear layer’s output dimension was half of its input dimension. We set the batch size to 300,000 to ensure the stability of the sampling distribution, thereby guaranteeing robust results. Additionally, the learning rate was initially set at 1e-4 and was polynomially decayed to 1e-8 within 10k epochs.
4 Emergence in ICL
@@
In this section, we focus on the performance of LLMs’ emergence in ICL, specifically examining the strength of emergence in the span of shots. Since existing datasets (e.g., SST-2 [25], AGNews [38], and EmoC [8]) do not meet the requirement of the same sequence length, we synthesized a set of simple few-shot samples having token length and positions aligned across different sequences. We curate three different datasets encompassing three different domains, each containing sequences of few-shot single-token entities with commas:
@@
  • Country: We select $25$ countries from the Vocabulary as entities, each represented by $1$ token (e.g., ‘Canada’, ‘Russia’). Each shot consists of one entity followed by a comma, totaling $2$ tokens. We constructed 25242322=303,6002524232230360025*24*23*22=303,60025 ∗ 24 ∗ 23 ∗ 22 = 303 , 600888The number of shots is decided to ensure the number of combinations in each category is over 300000. input sequences, each comprising $8$ tokens ($4$ different shots), such as “France, Mexico, Egypt, Russia,”.
@@
In the experiment, we observe that each entity, treated as a micro variable (i.e., the first token), produces similar mutual information across different positions. Consequently, in this section, we only use the entity in the first position to compute the mutual information to approximate the mean of all micro variables. Moreover, $E(l)$ also acts analogously in each block, so we utilize the mean of $\{E(l)\}^{L-1}_{l=0}$ to show the emergence strength. By varying the position $t$ of the macro variable (last token in a sequence), we use the following equation to compute the emergence strength at length $t$:
$\widehat{E}(t)=\frac{1}{L}\sum^{L-1}_{l=0}E(l)_{ma=t-1}=\frac{1}{L}\sum^{L-1}_{l=0}(MI(h^{t-1}_{l+1},h^{t-1}_{l})-MI(h^{0}_{l+1},h^{0}_{l}))$ (6)
4.1 Finding 1: Emergence Increases by Shots but Stabilizes within a Shot.
@@
Figure 2: $\widehat{E}(t)$ on 3 datasets with mean and variance.
Figure 2 indicates the emergence strength across the first 4 shots  [^9] The whole record of every token’s mutual information is detailed in Table 5, showing the example of GPT2-XL on the Animal category., based on which 3 observations can be derived:
  • The emergence escalates with an increasing number of shots within a certain range. (O1)
  • The standard deviation of emergence increases when the number of shots increases. (O2)
  • Tokens within the same shot exhibit approximately equivalent strength of emergence. (O3)
O1 highlights the role of “shot” in ICL: an increased number of shots enhances the confidence in predicting the next token. This observation corroborates substantial research that intuitively and experimentally explores ICL [19, 20, 32, 34, 15, 27]. Yet, the contribution of shots is not infinitely amplified. Table 3 indicates that the emergence tends to saturate beyond the $7$-th shot.
O2 indicates that emergence becomes unstable as the number of shots increases. To further study this observation, we explicitly report the emergence strength and standard deviation in Table 1 and compute the accuracy of the generations [^10] We randomly sample a total of $N=1000$ samples and regard the generation to be correct if the generated entity belongs to the corresponding domain of the dataset (country, animal or color) and is not repetitive. spanning over different number of shots. As can be seen from rows 1-9 of Table 1, as emergence strength ceases to grow and the variance reaches to the peak, the LM displays a higher probability of generating inaccurate responses. From a closer look, we discover that oftentimes, the LM fails to generate new entities due to “error repetition” (explanations and some examples can be found in Appendix B.3). This is aligned with existing study [37] related to hallucination. Specifically, LLMs struggle to correct itself subsequently after generating an erroneous output, consequently obstructing the emergence of correct new information, resulting in stagnation and fluctuation in emergence evaluation.
Emergence Statistics by each shot
@@
Table 1: The relationship between the accuracy of GPT2-XL outputs and emergence statistics by each shot in 3 categories.
However, this should not be confused with the power of ICL in exploiting more complex patterns effectively with more “shots” as the input. Different from the above observation, an increasing number of shots tends to bring higher accuracies under more complex scenarios. As shown in rows 10-15 of Table 1, we design four challenging tasks: ‘Asia’ and ‘Europe’ only provide countries in Asia and Europe, respectively, as input demonstrations; ‘Size’ contains animals arranged by size from smaller to larger; ‘Alphabet’ sorts the entities alphabetically based on the first letter. The accuracy results indicate that LLMs require more shots to capture complex patterns compared to simple patterns. Thus, it prompts us to conjecture if the stagnation and fluctuations in the emergence are associated with another hallucination: with excessive shots, LLMs may perceive more complex patterns beyond the surface (or even actual) appearance. In short, The correlation between emergence variance and hallucination would offer novel insights into the future development of hallucination detection and mitigation.
To validate O3, we compare the emergence strength ending with different positions (entity or comma) and vary the number of shots. Figure 4 demonstrates that the emergence strength remains consistent within each shot regardless of the position.
4.2 Finding 2: Emergence Is Associated with the Model, Dataset, Task, and Prompt.
To investigate the influence of different factors on emergence strength, we treat the performance of GPT2-XL on the “country” dataset as a baseline and implement a series of ablation studies. Appendix C details the variations when different factors change. It was observed that the emergence increases with the size of the model and the length of the token. This corroborates existing observational conclusions [31, 26] that emergence is more likely to occur in scenarios with a larger number of model parameters and richer contextual information. In addition, our study also identifies variations in emergence depend on the semantics, task, and prompt form, which also resonates with findings from prior research [18, 35, 17].
5 Emergence in Sentences and Datasets
The emergence of natural sentences is also a concern of our study. We randomly select 300,000 sequences, each consisting of $8$ tokens, from OpenOrca [16] and OpenHermes [29], respectively. OpenOrca and OpenHermes are both large-scale, multi-domain QA datasets. These sequences were selected to ensure that the first token in each sequence is the actual beginning of a sentence.
In our experiments, we observe potential discrepancies in mutual information for individual tokens at different positions within a sentence (i.e., micro mechanisms in different positions are not consistent). These discrepancies are detailed in Appendix D). Consequently, for the mutual information of micro-level variables, we adhere to Equation 3 which averages the micro MI at each position:
$\widehat{E}(t)=\frac{1}{L}\sum^{L-1}_{l=0}E(l)_{ma=t-1}=\frac{1}{L}\sum^{L-1}_{l=0}(MI(h^{t-1}_{l+1},h^{t-1}_{l})-\frac{1}{t}\sum^{t-1}_{m=0}MI(h^{m}_{l+1},h^{m}_{l}))$ (7)
5.1 Finding 3: Emergence Increases by Tokens in Natural Sentences
Figure 3 presents the performance of GPT2-large (0.8B), GPT2-XL (1B), GEMMA (2B), and OpenLlama (3B) on OpenOrca and OpenHermes datasets. We observe that emergence increases with the number of tokens, as natural sentences often possess sufficiently explicit semantics to provide confidence in predicting the next tokens. However, different datasets exhibit varying levels of emergence. Given the substantial size of our sample (300k), we attribute these variations primarily to differences in domain-specific characteristics. Moreover, we observed that the strength of emergence is averagely lower than that of ICL in Section 4.1. This highlights that ICL does enable LLMs to more effectively ascertain the next tokens compared to natural sentences.
(a) OpenOrca
(b) OpenHermes
Figure 3: $\widehat{E}(t)$ on 2 datasets.
Concurrently, disparate performances observed across the two datasets and three model families suggest that the domain of the training data and preprocessing methodologies are likely critical factors, as further supported by the evaluations of individual tokens at different sentence positions elaborated in Appendix D. For instance, GPT2-large and GPT2-XL, which share the same dataset (40GB WebText), preprocessing methods, and model architecture, exhibit a common characteristic: the mutual information of the first token is consistently lower than the average mutual information of micro-level variables.
5.2 Finding 4: Texts Generated from LLMs and Humans Exhibit Different Emergence
Motivated by the robust emergence in natural sentences in Section 5.1, we seek to measure the differences in text generated by larger language models compared to human texts, as well as the variations among these LLMs themselves. Specifically, we use questions from OpenHermes as inputs and collect responses by invoking the APIs of GPT-4, Claude3-opus, Claude3-sonnet, and Llama3. These responses were subjected to the same data processing methods described in Section 5. To evaluate these extremely large and closed-source language models, we implement a 3-step strategy:
Step 2: Following the data processing in Section 5, we format these answers into input sequences of $8$ tokens and obtained their representations using smaller LMs (e.g., GPT-2, GEMMA).
Step 3: These representations were then processed through an estimator to calculate the mutual information introduced in Section 3.2, thereby determining the emergence strengths of these answers via Equation 7.
Table 2 illustrates an interesting phenomenon: LLM-generated texts exhibited substantially greater emergence strength than human texts. This observation is intuitive—given that LLMs aim to generate tokens with the highest probability, naturally resulting in greater entropy reduction.
Another observation is that the text generated by different LLMs (GPT-4, Claude3, and Llama3) displays variations in emergence strengths. Significant differences are observed not only in the maximum strength of the emergence but also in the patterns of growth. Without actually computing the transformer representations of the target LLMs, these findings open a promising path towards estimating the emergence capability from extremely large and closed-source LMs without expensive computational costs.
@@ 5.2 Finding 4: Texts Generated from LLMs and Humans Exhibit Different Emergence
Table 2: Emergence in texts generated from human and popular LLMs. “text” refers to the party that generates the text. “Estimator” refers to the LM used to transform the text into representations and estimates the emergence strength using $f$ described in Section 3.2.

### v2

@@ 3.2 How to Estimate IE in LLMs?
It is not feasible to directly compute the mutual information in Eq. 3 using Kullback-Leibler (KL) divergence, as the input lies in a high-dimensional continuous space. To address that, we resort to an approximation method using mean values proposed in Belghazi et al. [6]:
@@
where $\theta$ denotes the parameters for $f$, ||||| | denotes the concatenation operation and $B$ is the batch size. $x,y\in\mathbb{R}^{D}$ are two inputs for computing the mutual information $MI(x,y)$. $x^{b}||y^{b}$ corresponds to sampling from the joint distribution $P_{XY}$, while $x^{b}||y^{i\neq b}$ corresponds to sampling from the marginal distribution $P_{X}$ and $P_{Y}$ [^4] In our implementation, the batch size was increased to encompass the entirety of the sample set to ensure the rationality of $P_{XY}$, $P_{X}$ and $P_{Y}$.. When $\mathcal{L}$ converges to the minimum $\hat{\mathcal{L}}$, we can obtain the final estimated mutual information as $MI(x,y)$=$-\log_{2}^{e}*\hat{\mathcal{L}}$. (More details and proofs are shown in Belghazi et al. [6].)
@@
To get the IE value $E(l)$ in Eq. 3, we compute $MI(h^{ma}_{l+1},h^{ma}_{l})$ by replacing $x^{b}$ and $y^{b}$ in Eq. 5 with $h^{ma}_{l+1,s}$ and $h^{ma}_{l,s}$ obtained by applying the LM to the same input sequence $s$, whereas replacing $y^{i\neq b}$ with $h^{ma}_{l,s^{\prime}}$ using a different sequence $s^{\prime}\neq s$. Similar operations apply when computing $MI(h^{mi\_t}_{l+1},h^{mi\_t}_{l})$. (Refer to Appendix A for a complete algorithm for estimator.)
4 Implementation
@@
Our algorithm requires that the number of samples is sufficiently large (over 300k) to provide a good estimate of the mutual information. Meanwhile, the length of each sequence within a dataset should be kept the same to facilitate position-wise observation and meaningful computation. Due to resource constraints, our comparative analysis is limited to GPT2-large (812M), GPT2-XL (1.61B), GEMMA (2.51B), and OpenLlama (3B) models. Fortunately, this parameter range is sufficient to observe variations and regularities of IE. For those LLMs with extremely large size or closed resource (e.g., GPT-4, Claude3, etc.), we design another efficient strategy that enables their IE evaluations as shown in Section 6.3. All computational experiments can be conducted on one NVIDIA GeForce RTX 3090 GPU. The estimator $f$ in Eq. 4 is a model of a 10-layer neural network comprising linear layers and leaky ReLU activation functions, where each linear layer’s output dimension was half of its input dimension. We set the batch size to 300,000, the learning rate to 1e-4, and polynomially decayed to 1e-8 within 10k epochs. We examine the IE value of LLMs under two distinct settings: ICL with few-shot examples and natural sentences without demonstrations.
4.1 ICL Scenario
@@
Since existing datasets (e.g., SST-2 [22], AGNews [32], and EmoC [8]) do not meet the requirement of the same sequence length, we synthesized a set of simple few-shot samples having token length and positions aligned across different sequences. We curate three different datasets encompassing three different domains, each containing sequences of few-shot single-token entities with commas:
@@
Country: We select $25$ countries from the Vocabulary as entities, each represented by $1$ token (e.g., ‘Canada’, ‘Russia’). Each shot consists of one entity followed by a comma, totaling $2$ tokens. We constructed 25242322=303,6002524232230360025*24*23*22=303,60025 ∗ 24 ∗ 23 ∗ 22 = 303 , 600555The number of shots is decided to ensure the number of combinations in each category is over 300000. input sequences, each comprising $8$ tokens ($4$ different shots), such as “France, Mexico, Egypt, Russia,”.
@@
In the experiment, we observe that each entity, treated as a micro variable (i.e., the first token), produces similar mutual information across different positions. Consequently, in this scenario, we only use the entity in the first position to compute the mutual information of micro variables (i.e., $\frac{1}{T}\sum^{T-1}_{t=0}MI(h^{mi\_t}_{l+1},h^{mi\_t}_{l})$ in Eqt 3 is replaced by $MI(h^{0}_{l+1},h^{0}_{l})$). Moreover, $E(l)$ also acts analogously in each block, so we utilize the mean of $\{E(l)\}^{L-1}_{l=0}$ to show the IE (i.e., $\widehat{E}(t)=\frac{1}{L}\sum^{L-1}_{l=0}E(l)$).
4.2 Natural Sentence Scenario
We randomly select 300,000 natural sequences, each consisting of $8$ tokens, from OpenOrca [14] and OpenHermes [26], respectively. OpenOrca and OpenHermes are both large-scale, multi-domain QA datasets. These sequences were selected to ensure that the first token in each sequence is the actual beginning of a sentence.
In our experiments, we observe potential discrepancies in mutual information for individual tokens at different positions within a sentence (i.e., micro mechanisms in different positions are not consistent). These discrepancies are detailed in Appendix C). Consequently, for the mutual information of micro-level variables, we keep the same as to Equation 3 which averages the micro MI at each position and utilizes the mean of each layer’s $E(l)$ to show the IE.
5 Informativeness
Figure 2: The increasement of IE, EM, Accuracy, and model loss for GPT2-XL in comparison to the previous token: $\text{increasement}=(value(t)-value(t-1))/value(t)$, where $value(t)$ represents the value at token $t$. Therefore, a positive increasement $(>0)$ indicates an increase in the metric value, and a decrease vice versa.
5.1 Semantics Faithfulness
We have designed several experiments to substantiate the semantics faithfulness of IE. We observed the change in IE and other popular metrics (Exact Match, Accuracy, and model loss) with the increase in the number of tokens, using the samples from the OpenOrca dataset. Figure 2 demonstrates the change in their values, with $\text{increasement}>0$ representing the value increasing from that in the previous token. Only IE consistently exhibits an upward trend (i.e., $>0$), which aligns with the intuition: what a sentence intends to convey is increasingly deterministic along with the increasing number of tokens. Moreover, the low variance (reflected as the shaded area in Figure 2) in IE values exhibits commendable stability compared to other metrics.
5.2 Semantics Sensitivity
Subsequently, we aim to examine the semantics sensitivity of IE, particularly its ability to reflect differences when minor perturbations are introduced into the semantics. Consequently, we conducted a series of ablation studies to modulate certain factors (such as dataset size, attributes, tasks, and format) individually. We treat the performance of GPT2-XL on the “country” dataset as a baseline. Appendix D details the variations when different factors are changed. It was observed that IE increases with the model’s size. This corroborates the rationale that a model with a larger size generally has better capability to determine semantics. In addition, our study also identifies variations in IE against different tasks and prompts, which also resonates with findings from prior research [16, 29, 15].
5.3 Connection to Emergence
(a) Task Performances
(b) IE values
Figure 3: IE and Model Performance with model size increasing in Arithmetic
Moreover, we demonstrate that IE manifests a steep ascend within the parameter range of $10^{8}$ to $10^{10}$ across 8 arithmetic tasks, which is detailed in Appendix B). Given the confines of computational resources, we were able to select 8 models within the parameter range of GPT2 ($1*10^{8}$), GPT2-large ($7*10^{8}$), GPT2-XL ($1*10^{9}$), Gemma ($2*10^{9}$), OpenLlama ($3*10^{9}$), GPT-J ($6*10^{9}$), Gemma ($7*10^{9}$), and GPT-NeoX ($2*10^{10}$). In light of the existing evaluation work, Big-Bench [23], we discovered the emergent phenomena within the arithmetic tasks emerge within the parameters of $[10^{8},10^{10}]$. Consequently, the association between the performance and IE values of 8 arithmetic tasks was investigated, as shown in Figure 3. For model performance, we directly adopt the default settings of the Big-Bench benchmark. As for IE, we took the average of the IE values of the initial five output tokens to be the final result.
A marked enhancement in task performance occurs once effective parameters reach $10^{10}$, thereby showcasing an emergent phenomenon. The average IE experiences a substantial surge within the same parameter range ($10^{9}$ to $10^{10}$). As a pioneering work proposing a quantitative metric to reflect the level of semantics deterministically, we believe our method could also greatly benefit further research on Emergence.
6 Findings
6.1 IE Increases Only when A New Demonstration Appears in ICL
@@
Figure 4: $\widehat{E}(t)$ on ICL and natural scenarios with mean and variance.
Figure 4 illustrates that IE naturally becomes higher with increased tokens. However, there is a strikingly different trend between ICL and natural scenarios (containing natural sentences). In a natural scenario, IE increases with each successive token and achieves a rapid convergence (around the 6th token), whereas, under the ICL scenario, IE only increases when a new demonstration emerges (at positions of the 2nd token, 4th token, 6th token), but with a higher upper bound and requiring more tokens to reach to the highest value.
We subsequently investigate how many demonstrations are needed before IE ceases to increase. Table 4 in the Appendix indicates that the three ICL categories under study tend to saturate at the 7th demonstration (though this does not suggests a generic ICL phenomenon). Moreover, we test whether increasing the number of tokens within each demonstration would maintain this "stepwise elevating" pattern. Figure 6 shows that the IE scores within each demonstration does not change when the length of each demonstration is increased to 5 tokens, 6 tokens, and 7 tokens.
Hence, we can interpret ICL’s role in enhancing semantics determinability: ICL bolsters semantics determinability via demonstrations, where increasing the number of demonstrations can increase the ability of capturing semantics beyond natural text, but eventually saturates after a certain quantity.
Concurrently, disparate performances observed across the two datasets and three model families suggest that the domain of the training data and preprocessing methodologies are likely critical factors, as further supported by the evaluations of individual tokens at different sentence positions elaborated in Appendix C. For instance, GPT2-large and GPT2-XL, which share the same dataset (40GB WebText), preprocessing methods, and model architecture, exhibit a common characteristic: the mutual information of the first token is consistently lower than the average mutual information of micro-level variables.
6.2 Higher IE with Large Standard Deviation Corresponds to Certain Hallucination
Figure 4666The complete record of every token’s mutual information is detailed in Table 5, showing the example of GPT2-XL on the Animal category. indicates that IE becomes unstable as the number of demonstrations increases. To further study this observation, we explicitly report the IE and standard deviation (s.d.) in Table 1 and compute the accuracy of the generations [^7] We randomly sample a total of $N=1000$ samples and regard the generation to be correct if the generated entity belongs to the corresponding domain of the dataset (country, animal or color) and is not repetitive. spanning over different numbers of shots. As can be seen from rows 1-9 of Table 1, as IE ceases to grow and the s.d. reaches peak, the LM displays a higher probability of generating inaccurate responses. From a closer look, we discover that oftentimes, the LM fails to generate new entities due to “error repetition” (explanations and some examples can be found in Appendix E.3). This is aligned with existing study [31] related to hallucination. Specifically, LLMs struggle to correct themselves after generating an erroneous output, resulting in stagnation and fluctuation in IE.
IE value by each shot
@@ 6.2 Higher IE with Large Standard Deviation Corresponds to Certain Hallucination
Table 1: The relationship between the accuracy of GPT2-XL outputs and IE by each shot in 3 categories.
@@
Table 2: IE in texts generated from human and popular LLMs. “text” refers to the party that generates the text. “Estimator” refers to the LM used to transform the text into representations and estimates the IE value using $f$ described in Section 3.2.
However, this should not be confused with the power of ICL in exploiting more complex patterns effectively with more “shots” as the input. Different from the above observation, an increasing number of shots tend to bring higher accuracies under more complex scenarios. As shown in rows 10-15 of Table 1, we design four challenging tasks: ‘Asia’ and ‘Europe’ only provide countries in Asia and Europe, respectively, as input demonstrations; ‘Size’ contains animals arranged by size from smaller to larger; ‘Alphabet’ sorts the entities alphabetically based on the first letter. The accuracy results indicate that LLMs require more shots to capture complex patterns compared to simple patterns. Thus, it prompts us to conjecture if the stagnation and fluctuations in the IE are associated with another hallucination: with excessive shots, LLMs may perceive more complex patterns beyond the surface (or even actual) appearance. In short, the correlation between IE s.d. and hallucination would offer novel insights into the future development of hallucination detection and mitigation.
6.3 Texts Generated from LLMs and Humans Exhibit Different IE Values
We seek to measure the differences in text generated by larger language models compared to human texts, as well as the variations among these LLMs themselves. Specifically, we use questions from OpenHermes as inputs and collect responses by invoking the APIs of GPT-4, Claude3-opus, Claude3-sonnet, and Llama3. These responses were subjected to the same data processing methods described in Section 4.2. To evaluate these extremely large and closed-source language models, we implement a 3-step strategy:
Step 2: Following the data processing in Section 4.2, we format these answers into input sequences of $8$ tokens and obtained their representations using smaller LMs (e.g., GPT-2, GEMMA).
Step 3: These representations were processed through an estimator to calculate the mutual information introduced in Section 3.2, thereby determining the IE values of these answers via Equation 3.
Table 2 illustrates an interesting phenomenon: LLM-generated texts exhibited substantially greater IE value than human texts. This observation is intuitive—given that LLMs aim to generate tokens with the highest probability, naturally resulting in greater entropy reduction.
Another observation is that the text generated by different LLMs (GPT-4, Claude3, and Llama3) displays variations in IE values. Significant differences are observed not only in the maximum strength of the IE but also in the patterns of growth. Without actually computing the transformer representations of the target LLMs, these findings open a promising path to estimate the semantics capturing capability from extremely large and closed-source LMs without expensive computational costs.
7 Conclusion
In this paper, we mathematically model the entropy of tokens and propose a quantitative metric, IE, representing the LLM’s ability to obtain semantics from tokens. Under the proposed low-resource estimator, we corroborate that IE possesses semantics faithfulness and sensitivity not found in other metrics. under the settings of ICL and natural sentences, We conducted extensive experiments explaining why ICL provides clearer semantics than natural text, as well as the intrinsic relationship between IE and hallucinations. Simultaneously, we discovered that IE can be utilized to distinguish whether the source of the text originates from humans or LLMs, particularly a simple and feasible strategy for those of significantly large LMs.